# Quintessential inflation studied through Semiclassical Methods

### v1

@@ 1 Introduction
In recent years, generalized $\alpha$–attractor inflationary models have attracted considerable attention because they have not only a well–understood origin in conformal and super–conformal field theories but also they have a close relationship with supergravity theories [22, 23, 24, 25, 26]. This means that there are also connections between these models and string theory, with some models being constructed using Kähler and super potentials that appear natural from a string theory perspective [27, 28].
Furthermore, their general predictions lead to a general prediction for inflationary observables and lie close to the center of current observational bounds on the primordial power spectra, namely the scalar spectral index $n_{S}(k)$ and the tensor–to–scalar ratio $r(k)$ which are consistent with observations from the Planck satellite [29, 30, 31].
The most recent exciting tension in cosmology is the Hubble tension, arising from a growing discrepancy between the $\Lambda CDM$–based predictions of the Hubble constant from cosmic microwave background measurements and direct measurements in the local universe. One of the most promising candidates to resolve this tension is the inclusion of an Early Dark Energy (EDE) component that can inject energy near the time of recombination, diluting quickly afterward to allow the EDE to smoothly transition into the observed late–time dark energy, thus alleviating the Hubble tension [32, 33, 34, 35]. $\alpha$–attractor inflationary models provide a unified theoretical framework that can naturally accommodate both EDE and late–time dark energy within the same scalar field potential [36, 37, 38].
@@ 2 The $\alpha$–attractor Inflationary Model
where $n$, and $\alpha$ are dimensionless parameters, and $n$ depends on $\alpha$ in the following way $n=\kappa\,\sqrt{6\alpha}$. The parameters $n$ and $\alpha$ are set for Saló in $n=124$ and $\alpha=10^{-2}$ [39]. For these parameters, the form of the $\alpha$–attractor potential Eq. (11) becomes, see Figure 1,
@@ 3.2.2 Fit of the Scalar Field $\varphi(t)$
In Table 1, we present the $\chi^{2}_{i}$ statistical factor for the slow–roll approximation and the fitted expression. From these results, we conclude that the fitted expression provides a better analytical description of the behavior of the scale factor $a(t)$ and the scalar field $\varphi(t)$ than the slow–roll approximation.
@@ 5.1 Numerical Integration
the integration is done from $t_{\mathrm{ini}}=0.001\,h_{c}(k)$ until $t_{\mathrm{end}}=0.05\,h_{c}(k)$, where $h_{c}(k)$ is the horizon defined by the time when the mode freezes $k=aH$. In this stage, we have used Eq. (33) as the initial condition but written in terms of $t$, that is $\tau\simeq-1/[a(t)H]$ since $H$ is nearly constant inside the horizon  [^1] In particular, during inflation, $\tau=\int^{a}_{0}{da}{Ha^{2}}\simeq{1}{H}\int^{a}_{0}{da}{a^{2}}\simeq-{1}{aH}.$ , then $u_{k}$ and $v_{k}$ exhibit oscillatory behavior.
@@ 5.3 Uniform approximation method
$\ddot{U}_{k}+R_{\mathrm{S}}(k,t)U_{k}$ $=$ $0,$ (53)
$\ddot{V}_{k}+R_{\mathrm{T}}(k,t)V_{k}$ $=$ $0,$ (54)
with,
where $U(k)$ satisfies the asymptotic conditions
the asymptotic conditions (57) and (58) also hold for $V_{k}$.
For the computation of the power spectrum we need to take the limit $k\,t\rightarrow\infty$ of the solutions (8(b)) and (8(b)). In this limit we have
Finally,
where $\mathrm{ua}$’ means the uniform approximation method.
@@ 5.4 Phase–integral method
Let us consider the differential equation,
$\dfrac{\mathrm{d}^{2}u_{k}}{\mathrm{d}z^{2}}+R(z)\,u_{k}=0,$ (84)
$u_{k}=q^{-\nicefrac{{1}}{{2}}}(z)\exp\left[\pm i\,\omega(z)\right],$ (85)
where
$\omega(z)=\int^{z}q(z)\,\mathrm{d}z.$ (86)
$\epsilon_{0}=Q^{-\nicefrac{{3}}{{@}}}(z)\dfrac{\mathrm{d}^{2}}{\mathrm{d}z^{2}}Q^{-\nicefrac{{1}}{{2}}}(z)+\dfrac{R(z)-Q^{2}(z)}{Q^{2}(z)},$ (88)
$\xi=\int^{z}Q(z)\,\mathrm{d}z,$ (91)
we obtain,
$\dfrac{q\lambda}{Q}=\sum^{\infty}_{n=0}Y_{2n}\lambda^{2n}.$ (93)
Substituting Eq. (93) in Eq. (92), we obtain,
$Y_{0}$ $=$ $1,$ (95)
$Y_{2}$ $=$ $\dfrac{1}{2}\,\epsilon_{0},$ (96)
$Y_{4}$ $=$ $-\dfrac{1}{8}\left(\epsilon_{0}^{2}+\epsilon_{2}\right),$ (97)
where $\epsilon_{\nu}$ is defined as,
Truncating the series (93) at $n=N$ with $\lambda=1$ we obtain
$q(z)=Q(z)\sum^{N}_{n=0}Y_{2n},$ (101)
Substituting Eq. (101) into Eq. (86) we have,
$\omega(z)=\sum_{n=0}^{N}\omega_{2n}(z),$ (102)
where,
$\omega_{2n}(z)=\int^{z}Y_{2n}Q(z)\,\mathrm{d}z.$ (103)
$\omega_{2n}(z)=\dfrac{1}{2}\int_{\Gamma_{t}}Y_{2n}(z)Q(z)\,\mathrm{d}z,$ (104)
and for $t>t_{\mathrm{ret}}$
where
$f_{2_{\mathrm{S}}}(k,t)$ $=$ $Y_{2_{\mathrm{S}}}(k,t)Q_{\mathrm{S}}(k,t),$ (117)
$f_{2_{\mathrm{T}}}(k,t)$ $=$ $Y_{2_{\mathrm{T}}}(k,t)Q_{\mathrm{T}}(k,t).$ (118)
$f_{2_{\mathrm{S}}}(k,t)$ $=$ $A(k,t)(t-\tau_{\mathrm{S}})^{\nicefrac{{-5}}{{2}}},$ (119)
$f_{2_{\mathrm{T}}}(k,t)$ $=$ $B(k,t)(t-\tau_{\mathrm{T}})^{\nicefrac{{-5}}{{2}}},$ (120)
$c_{1}$ $=$ $-\mathrm{i}\,c_{2},$ (121)
$d_{1}$ $=$ $-\mathrm{i}\,d_{2},$ (123)
where $\mathrm{pi}$’ means the third–order phase integral method.
@@ 8 Conclusions
In this work, we solved numerically the scalar and tensor perturbation equations for the $\alpha$–attractor inflationary model using the slow–roll approximation, the uniform approximation method, and the third–order phase–integral method. This inflationary model allows us to describe both the initial accelerated expansion of the Universe during the inflationary epoch and the current accelerated expansion with the same potential. Once the scalar power spectrum $P_{\mathrm{S}}(k)$, the tensor power spectrum $P_{\mathrm{T}}(k)$, and the observables are calculated, we find that the third–order phase–integral method provides excellent agreement with the numerical results, providing more accurate results than both the slow–roll and uniform approximation methods. To compare the power spectra obtained from each approximation method with the numerical results, we employ the statistical factor $\chi_{i}^{2}$, and to compare the observables calculated with each approximation method to those from the numerical solution, we use the percentage relative error.

### v2

@@ 1 Introduction
In recent years, generalized $\alpha$–attractor inflationary models have attracted considerable attention because they have not only a well–understood origin in conformal and super–conformal field theories but also they have a close relationship with supergravity theories [22, 23, 24, 25, 26]. This means that there are also connections between these models and string theory, with some models being constructed using Kähler and super potentials that appear natural from a string theory perspective [27, 28]. Furthermore, their general predictions lead to a general prediction for inflationary observables and lie close to the center of current observational bounds on the primordial power spectra, namely the scalar spectral index $n_{S}(k)$ and the tensor–to–scalar ratio $r(k)$ which are consistent with observations from the Planck satellite [29, 30, 31].
The most recent exciting tension in cosmology is the Hubble tension, arising from a growing discrepancy between the $\Lambda CDM$–based predictions of the Hubble constant from cosmic microwave background measurements and direct measurements in the local universe. One of the most promising candidates to resolve this tension is the inclusion of an Early Dark Energy (EDE) component that can inject energy near the time of recombination, diluting quickly afterward to allow the EDE to smoothly transition into the observed late–time dark energy, thus alleviating the Hubble tension [32, 33, 34, 35]. The $\alpha$–attractor inflationary models provide a unified theoretical framework that can naturally accommodate both EDE and late–time dark energy within the same scalar field potential [36, 37, 38].
@@ 2 The $\alpha$–attractor Inflationary Model
where $n$, and $\alpha$ are dimensionless parameters, and $n$ depends on $\alpha$ in the following way $n=\kappa\,\sqrt{6\alpha}$. The parameters $n$ and $\alpha$ are set by the authors in Saló in $n=124$ and $\alpha=10^{-2}$ [39]. For these parameters, the form of the $\alpha$–attractor potential Eq. (11) becomes, see Figure 1,
@@ 3.2.2 Fit of the Scalar Field $\varphi(t)$
In Table 1, we present the $\chi^{2}_{i}$ statistical factor for the slow–roll approximation and the fitted expression. From these results, we conclude that the fitted expression provides a better analytical description of the behavior of the scale factor $a(t)$ and the scalar field $\varphi(t)$ than the slow–roll approximation. It is important to notice the relatively high value of $\chi^{2}_{\varphi}$, in comparison with $\chi^{2}_{a}$, this is because the value of $\varphi(t)$ with each approximate method deviates more from the numerical result that the scale factor $a(t)$.
@@ 5.1 Numerical Integration
the integration is done from $t_{\mathrm{ini}}=0.001\,h_{c}(k)$ until $t_{\mathrm{end}}=0.05\,h_{c}(k)$, where $h_{c}(k)$ is the horizon defined by the time when the mode freezes $k=aH$. In this stage, we have used Eq. (33) as the initial condition but written in terms of $t$, that is $\tau\simeq-1/[a(t)H]$ since $H$ is nearly constant inside the horizon, then $u_{k}$ and $v_{k}$ exhibit oscillatory behavior.
@@ 5.3 Uniform approximation method
This method is discussed in the Appendix.
@@ 5.4 Phase–integral method
This method is discussed in the Appendix.
@@ 8 Conclusions
In this work, we solved numerically the scalar and tensor perturbation equations for the $\alpha$–attractor inflationary model using the slow–roll approximation, the uniform approximation method, and the third–order phase–integral method. This inflationary model allows us to describe both the initial accelerated expansion of the Universe during the inflationary epoch and the current accelerated expansion with the same potential. Once the scalar power spectrum $P_{\mathrm{S}}(k)$, the tensor power spectrum $P_{\mathrm{T}}(k)$, and the observables are calculated, we find that the third–order phase–integral method provides excellent agreement with the numerical results, providing more accurate results than both the slow–roll and uniform approximation methods, at least for the scalar power spectrum. However the slow–roll approximation performs best results for the tensor power spectrum. Which can be deduced from the comparison of the power spectra obtained from each approximation method with the numerical results, for which we employ the statistical factor $\chi_{i}^{2}$. Also we compare the observables of the $\alpha$–attractor inflationary model calculated with each approximation method to those obtained from the numerical solution, using the percentage relative error, obtained better results with those calculated with the $3^{\textnormal{rd}}$–order phase–integral–method. Additionally we compare our model with the Starobinsky inflationary model and we can conclude that both models are a good representation of the inflationary epoch because both lie within the $95\%$ confidence region.
9 Appendix
9.1 Uniform approximation method
$\ddot{U}_{k}+R_{\mathrm{S}}(k,t)U_{k}$ $=$ $0,$ (65)
$\ddot{V}_{k}+R_{\mathrm{T}}(k,t)V_{k}$ $=$ $0,$ (66)
with,
where $U(k)$ satisfies the asymptotic conditions
the asymptotic conditions (69) and (70) also hold for $V_{k}$.
For the computation of the power spectrum we need to take the limit $k\,t\rightarrow\infty$ of the solutions (11(b)) and (11(b)). In this limit we have
Finally,
where $\mathrm{ua}$’ means the uniform approximation method.
9.2 Phase–integral method
Let us consider the differential equation,
$\dfrac{\mathrm{d}^{2}u_{k}}{\mathrm{d}z^{2}}+R(z)\,u_{k}=0,$ (96)
$u_{k}=q^{-\nicefrac{{1}}{{2}}}(z)\exp\left[\pm i\,\omega(z)\right],$ (97)
where
$\omega(z)=\int^{z}q(z)\,\mathrm{d}z.$ (98)
$\epsilon_{0}=Q^{-\nicefrac{{3}}{{2}}}(z)\dfrac{\mathrm{d}^{2}}{\mathrm{d}z^{2}}Q^{-\nicefrac{{1}}{{2}}}(z)+\dfrac{R(z)-Q^{2}(z)}{Q^{2}(z)},$ (100)
$\xi=\int^{z}Q(z)\,\mathrm{d}z,$ (103)
we obtain,
$\dfrac{q\lambda}{Q}=\sum^{\infty}_{n=0}Y_{2n}\lambda^{2n}.$ (105)
Substituting Eq. (105) in Eq. (104), we obtain,
$Y_{0}$ $=$ $1,$ (107)
$Y_{2}$ $=$ $\dfrac{1}{2}\,\epsilon_{0},$ (108)
$Y_{4}$ $=$ $-\dfrac{1}{8}\left(\epsilon_{0}^{2}+\epsilon_{2}\right),$ (109)
where $\epsilon_{\nu}$ is defined as,
Truncating the series (105) at $n=N$ with $\lambda=1$ we obtain
$q(z)=Q(z)\sum^{N}_{n=0}Y_{2n},$ (113)
Substituting Eq. (113) into Eq. (98) we have,
$\omega(z)=\sum_{n=0}^{N}\omega_{2n}(z),$ (114)
where,
$\omega_{2n}(z)=\int^{z}Y_{2n}Q(z)\,\mathrm{d}z.$ (115)
$\omega_{2n}(z)=\dfrac{1}{2}\int_{\Gamma_{t}}Y_{2n}(z)Q(z)\,\mathrm{d}z,$ (116)
and for $t>t_{\mathrm{ret}}$
where
$f_{2_{\mathrm{S}}}(k,t)$ $=$ $Y_{2_{\mathrm{S}}}(k,t)Q_{\mathrm{S}}(k,t),$ (129)
$f_{2_{\mathrm{T}}}(k,t)$ $=$ $Y_{2_{\mathrm{T}}}(k,t)Q_{\mathrm{T}}(k,t).$ (130)
$f_{2_{\mathrm{S}}}(k,t)$ $=$ $A(k,t)(t-\tau_{\mathrm{S}})^{\nicefrac{{-5}}{{2}}},$ (131)
$f_{2_{\mathrm{T}}}(k,t)$ $=$ $B(k,t)(t-\tau_{\mathrm{T}})^{\nicefrac{{-5}}{{2}}},$ (132)
$c_{1}$ $=$ $-\mathrm{i}\,c_{2},$ (133)
$d_{1}$ $=$ $-\mathrm{i}\,d_{2},$ (135)
where $\mathrm{pi}$’ means the third–order phase integral method.